A bartender mixes a batch of vodka (40% alcohol) with tonic water (no alcohol) to make 5 liters of
vodka and tonic that is 22% alcohol. How much tonic water, and how much vodka does he need to
mix?

Answer by stanbon(75887)

A bartender mixes a batch of vodka (40% alcohol) with tonic water (no alcohol) to make 5 liters of vodka and tonic that is 22% alcohol.
How much tonic water, and how much vodka does he need to mix?
----------------------
0.4v + 0*t = 0.22*5
----
v = 1.1/0.4
v = 2.75 liters (amt of vodka needed)
----
5-2.75 = 1.25 liters (amt of tonic water needed)
---------------
Cheers,
Stan H.
--------------